You are given the following information.

Savings S = 150
Investment I = 100
Taxes T = 250
Government Purchases G = 500

Compute the level of private savings, public savings, and national savings.
1
Expert's answer
2020-09-17T08:38:27-0400

Answers

Private sector savings = $150

public sector savings = -$250

National savings = -$100


Solutions

Private savings are the private sector savings. That is, savings by the households and private firms.


Private savings = Y - ( T + C)

= Yd - C

= S

= $150



Public savings are the government savings. These savings represent the excess of government expenditure (G) over tax revenue (T).

Thus,

Public savings = T - G

= $250 - $500

= -$250


The public savings is negative. This means that government expenditure exceeded revenue (tax revenue) hence a deficit. Thus, there is a government budget deficit of $250.



National savings combine the private sector savings and the public sector savings.

National savings

= private savings + public savings

= $150 + (-$250)

= $150 - $250

= -$100


The national savings is negative meaning that the nation as a whole has a deficit; national spending exceeded national income. This negatively impacts money in circulation, availability of loanable funds, and economic prosperity as it represents a withdrawal from the circular flow of national income.